# Video[Video][video]

### v2

@@ 2 Background and prior work
We start by describing a framework for reconstructing material properties from microscopic imaging data. By representing material properties and data as vectors, we describe our reconstructions as the solution to an optimization problem. Using broadly applicable assumptions of linearity, shift invariance, and weak scattering, we describe all contrasts as the result of banks of point spread functions. We describe the physics behind the major contrast mechanisms and compute their transfer functions from illumination, scattering, and detection models. Throughout, we use carefully chosen basis functions to enable the physical interpretation of light and material properties. We describe all notation as we introduce it, and we collect all symbols in Supplement 1.
3.1 Objects and data as vectors, imaging as an operator
@@
We represent a biological sample as a series of volumetric maps of material properties, shown schematically in Figure 1a, i. For example, a cell might be approximately described by two fluorophore density maps, $f_{1}(\mathbf{r}_{o})$ and $f_{2}(\mathbf{r}_{o})$ where $\mathbf{r}_{o}$ is a 3D object position vector, one map for each of two different fluorophores that label biological structures of interest, and a density map $\rho(\mathbf{r}_{o})$. Three maps are unlikely to describe the sample completely, so we generalize and collect any number of volumetric maps into a single vector
@@
Figure 1: A framework for volumetric microscopy of phase, absorption, birefringence, diattenuation, fluorescence density, and fluorescence orientation. (a) Microscopes encode (i) true material properties into (ii) multi-contrast multi-channel volumetric datasets by changing spectral and polarization illumination and detection filters. Physics-informed models of the encoding process, $\mathcal{H}$, lead to physics-informed reconstruction operators, $\mathcal{R}$, that enable (iii) reconstructions of underlying material properties. (iv) We find that physics-informed reconstructions are useful for improving performance on downstream observation, biophysical, and learned tasks. (b) We model the encoding process of an (i) off-the-shelf inverted microscope stand outfitted with (ii) an incoherent source, variable spectral and polarization filters, and a polarization-sensitive detector. Two classes of contrast are available: label-free contrast where (iii) direct fields scatter coherently from the sample and direct-scatter interference generates contrast, and (iv) fluorescence contrast where direct fields that excite inelastic-scattering fluorophores are filtered so that scatter-scatter interference generates contrast. We model coherent scatterers as (v) a volume of scattering potential tensors that model the relationship between incident electric fields and induced dipoles moments. We model inelastic fluorescence scatterers as (vi) a volume of dipole emission moments. Both coherent and incoherent dipoles can be linear, circular, and elliptical, though fluorescent dipoles used in biological microscopy are primarily linear. (c) Expanding the scattering potential tensor onto spherical harmonic tensors enables interpretation in terms of phase, absorption, birefringence, and diattenuation. (d) Expanding dipole emitters into Cartesian components with complex coefficients enables interpretation of single emitters’ orientations. (e) Expanding measurable data into Stokes parameters enables modeling of all possible polarization measurements. See also Video 1.
When we image our sample in a microscope, we arrange for the material properties to be encoded into a list of volumetric datasets, each called a channel, shown schematically in Figure 1a, ii. For example, we might use a fluorescence light path to encode fluorophore density maps into two channels, $d_{1}(\mathbf{r}_{d})$ and $d_{2}(\mathbf{r}_{d})$ where $\mathbf{r}_{d}$ is a 3D detector position vector, then we can change to a transmission light path to encode the density $\rho(\mathbf{r}_{o})$ into the third channel $d_{3}(\mathbf{r}_{d})$. Similar to our object properties, we collect any number of channels into a single vector
@@
We can represent the imaging process with a single forward operator $\mathcal{H}$ that encodes the material properties $\mathbf{f}$ into measured volumetric datasets $\mathbf{d}$
@@
3.2 Reconstructing object properties
@@
$Q_{\eta}^{\text{(tls)}}(\mathbf{f},\mathbf{d})=\left\|\mathbf{d}-\mathbf{b}-\mathcal{H}\mathbf{f}\right\|_{2}^{2}-\eta\|\mathbf{f}\|_{2}^{2},$ (5)
@@
After choosing an objective function, the minimization problem needs to be solved, an often challenging task. Fortunately, if $\mathcal{H}$ is linear then the Tikhonov-regularized least-squares objective can be minimized in a single step described in Supplement 5, generating efficient noise-tolerant estimates of material properties, Figure 1a, iii.
As illustrated in Figure 1a, iv, physics-informed reconstruction operators prepare raw microscopy data for improved direct visual inspection, improved biophysical estimates [1], and improved performance on tasks performed by trained networks [5]. Therefore, the next sections develop physically interpretable models of light-matter interactions for diverse contrast modes.
3.3 Contrast modes
We model the image formation process of an off-the-shelf inverted microscope stand (Figure 1b, i–ii) outfitted with a spatially incoherent Köhler source with variable polarization filters, variable spectral filters, a high-NA condenser and objective, and an imaging path that includes a polarization-sensitive detector.
All contrast is formed by illuminating the sample with electric fields that scatter from the sample then interfere on the detector. If we consider only single scattering events, the first Born approximation, then we can rewrite Equation 3 as
@@ 3.3 Contrast modes
We consider two classes of contrast. Label-free contrast (Figure 1b, iii) is generated by illuminating the sample with light that interacts with the sample coherently—that is, scattered fields have the same wavelength and a fixed phase relationship with the illuminating fields. When a plane wave encounters a coherent scatterer, the oscillating electric field accelerates bound electrons in the scatterer, and these accelerated charges generate spherical scattered fields. The direct and scattered fields interfere and generate contrast via the scatter-direct and direct-scatter terms. The direct-direct term creates a uniform background, and for weakly scattering samples the scatter-scatter term is small and ignorable. Therefore, label-free contrast is generated by the direct-scatter and scatter-direct terms on top of a direct-direct background. Finally, each point on the source emits incoherently, so we can treat each source point individually and find the complete contrast pattern by summing over the source.
@@
Fluorescence contrast (Figure 1b, iv) is generated by illuminating fluorescent scatterers and imaging their scattered light. Fluorescent scatterers are incoherent, so the scattered fields have a random phase at a longer wavelength than the illuminating fields. Therefore, the scatter-direct and direct-scatter terms do not generate contrast, so the only way to measure sample-dependent contrast is via the small scatter-scatter term. Fortunately, the direct and scattered fields are at different wavelengths, so the direct fields can be filtered with minimal bleedthrough. Therefore, fluorescence contrast is generated by the scatter-scatter term with a direct-direct bleedthrough background. Finally, fluorescent scatterers emit incoherently, so we can find the complete contrast pattern by summing over the sample.
@@
Both label-free and fluorescence contrast modes can generate additional contrast from anisotropic samples. Label-free samples can be anisotropic if the scatterer’s bound electrons accelerate anisotropically. We illustrate a label-free anisotropic sample schematically as an electron bound to its nucleus by springs of varying spring constant (Figure 1b, v). When polarized light is incident on an anisotropic sample, it accelerates the bound electrons in linear, circular, or elliptical dipoles, which emit anisotropic polarized light in patterns that encode the orientation of the induced electron motion and the underlying anisotropy of the scatterer. Therefore, information about the sample’s label-free anisotropy is encoded in the polarization and intensity pattern of the detected light. Similarly, fluorescent scatterers emit along linear, circular, or elliptical dipoles (Figure 1b, vi), though linear dipoles are most common among the fluorophores used in biological microscopy.
3.4 Physically interpretable basis functions
@@
When we illuminate a label-free sample, the 3D induced dipole moment is the product of the incident field and a 3 $\times$ 3 matrix called the permittivity tensor [29, 21]. By convention, we change to a unitless quantity and subtract the isotropic background (Supplement 6) to arrive at a complete set of label-free sample properties—the complex-valued 3 $\times$ 3 matrix called the scattering potential tensor, $f^{(\text{lf})}_{ij}$. Each entry of the scattering potential tensor can be interpreted directly (e.g. the complex-valued $f^{(\text{lf})}_{xz}$ is the relative magnitude and phase of the $x$ component of the dipole induced by a $z$-oriented field), but this interpretation can be challenging to understand physically. To improve physical interpretability, we expand the scattering potential tensor onto the spherical harmonic tensors, a set of nine 3 $\times$ 3 matrices whose complex-valued expansion coefficients can be directly interpreted in terms of phase, absorption, birefringence, and diattenuation. We schematize each of these spherical harmonic tensors in Figure 1c by drawing each tensor’s eigenvalues and eigenvectors, and we describe the spherical harmonic tensor basis in detail in Supplement 6.
@@
In a fluorescent sample, the 3D emission dipole moment can be represented by a three-component vector $f^{(\text{fl})}_{i}$ (Figure 1d) with real-valued coefficients for purely linear dipoles and complex-valued coefficients for arbitrary dipoles. Contrast arises from the scatter-scatter term, so our measurements are proportional to the squares of the dipole components $|f^{(\text{fl})}_{i}|^{2}$.
@@
For dynamic ensembles of fluorescent emitters, the measurements are proportional to the second-moment matrix $\langle f^{(\text{fl})}_{i}f^{*(\text{fl})}_{j}\rangle$ [13, 34, 23]. Similar to the scattering potential tensor, we can expand the second-moment matrix onto the spherical harmonic tensors, but here we interpret the coefficients in terms of orientation distribution functions [14, 24].
@@
Finally, we express our data in terms of the Stokes parameters, a set of four real-valued parameters that are physically interpretable as the intensities measured behind various polarizing filters. The Stokes parameters can also be interpreted as the coefficients of the electric field’s second-moment matrix expanded onto the Pauli matrices (Figure 1e, Supplement 7).
@@
3.5 Linear contrast-separable shift-invariant imaging systems
Under approximations that are applicable to a wide range of microscopes (details in Section 5 and Supplement 8), we can model our imaging system by splitting our material properties, data channels, and forward model into independent groups called contrast modes, e.g. one label-free mode and several fluorescence modes. Each of these contrast modes is approximately linear across object properties, and each property-to-channel mapping is approximately spatially linear and shift invariant. Therefore, we can express our imaging model as
[…]
@@ 3.5 Linear contrast-separable shift-invariant imaging systems
where $m$ indexes contrast $m$odes, $p$ indexes material $p$roperties, $c$ indexes data $c$hannels, and $h^{(m)}_{cp}$ is a bank of point spread functions that model the entire multi-contrast multi-channel imaging system. We can reexpress this relationship in the Fourier domain as
@@
3.6 Summary of transfer functions
@@
The Green’s tensor spectrum is particularly important for modeling anisotropic contrast. Linear dipole moments emit polarized light in a doughnut-shaped intensity pattern (Figure 2a), and the Green’s tensor spectrum (Figure 2b) efficiently models all dipole emitters (coherent or incoherent; linear, circular, or elliptical dipoles in any orientation) with a single function.
Figure 2: Key components of the forward model. (a) A dipole emitter (blue arrow) emits polarized fields (red arrows) in an anisotropic pattern (black line, radius proportional to power). (b) Arbitrary linear, circular, and elliptical dipole emissions can be modeled with the Green’s tensor spectrum $G_{ij}(\boldsymbol{\nu})$. Each surface shows the quarter-maximum intensity of on-shell spectral field components (rows) $e_{x}$, $e_{y}$, and $e_{z}$ emitted by dipoles oriented along (columns) $\hat{\textbf{x}}$, $\hat{\textbf{y}}$, and $\hat{\textbf{z}}$, with relative phase shown in color (see color rose). Orthogonal slices through the Green’s tensor spectrum and the Green’s tensor are shown in Figure S3. (c–f) 3D support of various transfer functions with phase encoded in color. (c-d) Real label-free vector transfer functions with $\text{NA}_{\text{det}}=0.75$ and circularly polarized illumination that expresses polarization-resolved data (rows) as outputs of filters that modulate material properties (columns) with (c) low-NA illumination $\text{NA}_{\text{ill}}=0.1$ and (d) high-NA illumination $\text{NA}_{\text{ill}}=0.5$. (e) Scalar label-free transfer function with $\text{NA}_{\text{ill}}=0.5$ and $\text{NA}_{\text{det}}=0.75$. (f) Scalar fluorescence transfer function with $\text{NA}_{\text{det}}=0.75$. See also Video 2.
All transfer functions can be expressed as products and autocorrelations of the illumination, scattering, and detection models, see Table 1 and Supplement 8. We refer to the complete transfer functions as vector models because they account for the complete vectorial nature of light and dipole scattering. We also include scalar models that ignore vector effects, which are reasonable approximations when unpolarized illumination and unpolarized detection are used on isotropic samples.
vector scalar
Table 1: Transfer functions for different contrast modes (rows) and optical models (columns) expressed in terms of the illumination model $S$, the scattering model $G$, the detection model $P$, the object-space basis $\mathcal{Y}$, and the data-space basis $\sigma$. Sums over $i,i^{\prime},j,j^{\prime},k$ and $k^{\prime}$ are implied.
Figures 2c–f show the support and phase of several examples of waveOrder’s transfer functions. We briefly highlight several key features

### v3

@@ 4.2 WaveOrder framework
We start by describing a framework for reconstructing material properties from microscopic imaging data. By representing material properties and data as vectors, we describe our reconstructions as the solution to an optimization problem. Using broadly applicable assumptions of linearity, shift invariance, and weak scattering, we describe all contrasts as the result of banks of point spread functions. We describe the physics behind the major contrast mechanisms and compute their transfer functions from illumination, scattering, and detection models. Throughout, we use carefully chosen basis functions to enable the physical interpretation of light and material properties.
We describe all notation as we introduce it, and we collect all symbols in Supp. 1.
4.2.1 Objects and data as vectors
@@
We represent a biological sample as a series of volumetric maps of material properties, shown schematically in Fig. 1c, i. For example, a cell might be approximately described by two fluorophore density maps, $f_{1}(\mathbf{r}_{o})$ and $f_{2}(\mathbf{r}_{o})$ where $\mathbf{r}_{o}$ is a 3D object position vector, one map for each of two different fluorophores that label biological structures of interest, and a density map $\rho(\mathbf{r}_{o})$. Three maps are unlikely to describe the sample completely, so we generalize and collect any number of volumetric maps into a single vector
@@
When we image our sample in a microscope, we arrange for the material properties to be encoded into a list of volumetric datasets, each called a channel, shown schematically in Fig. 1c, ii. For example, we might use a fluorescence light path to encode fluorophore density maps into two channels, $d_{1}(\mathbf{r}_{d})$ and $d_{2}(\mathbf{r}_{d})$ where $\mathbf{r}_{d}$ is a 3D detector position vector, then we can change to a transmission light path to encode the density $\rho(\mathbf{r}_{o})$ into the third channel $d_{3}(\mathbf{r}_{d})$. Similar to our object properties, we collect any number of channels into a single vector
@@
4.2.2 Imaging and reconstruction as linear operators
We can represent the imaging process with a single forward operator $\mathcal{H}$ that encodes material properties $\mathbf{f}$ into measured volumetric datasets $\mathbf{d}$
@@ 4.2.2 Imaging and reconstruction as linear operators
4.2.3 Reconstructing object properties
@@
$Q_{\eta}^{\text{(tls)}}(\mathbf{f},\mathbf{d})=\left\|\mathbf{d}-\mathbf{b}-\mathcal{H}\mathbf{f}\right\|_{2}^{2}+\eta\|\mathbf{f}\|_{2}^{2},$ (5)
@@
After choosing an objective function, the minimization problem needs to be solved, an often challenging task. Fortunately, if $\mathcal{H}$ is linear then the Tikhonov-regularized least-squares objective can be minimized in a single step (Methods 4.2.9), generating efficient noise-tolerant estimates of material properties (Fig. 1c, iii).
4.2.4 Label-free and fluorescence contrast in a unified framework
All contrast is formed by illuminating the sample with electric fields that scatter from the sample then interfere on the detector. If we consider only single scattering events, the first Born approximation, then we can rewrite Eq. 3 as
@@ 4.2.4 Label-free and fluorescence contrast in a unified framework
We consider two classes of contrast. Label-free contrast (Fig. 1a, iii) is generated by illuminating the sample with light that interacts with the sample coherently—that is, scattered fields have the same wavelength and a fixed phase relationship with the illuminating fields. When a plane wave encounters a coherent scatterer, the oscillating electric field accelerates bound electrons in the scatterer, and these accelerated charges generate spherical scattered fields. The direct and scattered fields interfere and generate contrast via the scatter-direct and direct-scatter terms. The direct-direct term creates a uniform background, and for weakly scattering samples the scatter-scatter term is small and ignorable. Therefore, label-free contrast is generated by the direct-scatter and scatter-direct terms on top of a direct-direct background. Finally, each point on the source emits incoherently, so we can treat each source point individually and find the complete contrast pattern by summing over the source.
@@
Fluorescence contrast (Fig. 1a, iv) is generated by illuminating fluorescent scatterers and imaging their scattered light. Fluorescent scatterers are incoherent, so the scattered fields have a random phase at a longer wavelength than the illuminating fields. Therefore, the scatter-direct and direct-scatter terms do not generate contrast, so the only way to measure sample-dependent contrast is via the small scatter-scatter term. Fortunately, the direct and scattered fields are at different wavelengths, so the direct fields can be filtered with minimal bleedthrough. Therefore, fluorescence contrast is generated by the scatter-scatter term with a direct-direct bleedthrough background. Finally, fluorescent scatterers emit incoherently, so we can find the complete contrast pattern by summing over the sample.
@@
Both label-free and fluorescence contrast modes can generate additional contrast from anisotropic samples. Label-free samples can be anisotropic if the scatterer’s bound electrons accelerate anisotropically. We illustrate a label-free anisotropic sample schematically as an electron bound to its nucleus by springs of varying spring constant (Fig. 1a, v). When polarized light is incident on an anisotropic sample, it accelerates the bound electrons in linear, circular, or elliptical dipoles, which emit anisotropic polarized light in patterns that encode the orientation of the induced electron motion and the underlying anisotropy of the scatterer. Therefore, information about the sample’s label-free anisotropy is encoded in the polarization and intensity pattern of the detected light. Similarly, fluorescent scatterers emit along linear, circular, or elliptical dipoles (Fig. 1a, vi), though linear dipoles are most common among the fluorophores used in biological microscopy.
4.2.5 Physically interpretable basis functions
@@
When we illuminate a label-free sample, the 3D induced dipole moment is the product of the incident field and a $3\times 3$ matrix called the permittivity tensor [yeh_permittivity_2024, wolf_introduction_2007]. By convention, we change to a unitless quantity and subtract the isotropic background (Supp. 3) to arrive at a complete set of label-free sample properties—the complex-valued 3 $\times$ 3 matrix called the scattering potential tensor, $f^{(\text{lf})}_{ij}$. Each entry of the scattering potential tensor can be interpreted directly (e.g. the complex-valued $f^{(\text{lf})}_{xz}$ is the relative magnitude and phase of the $x$ component of the dipole induced by a $z$-oriented field), but this interpretation can be challenging to understand physically. To improve physical interpretability, we expand the scattering potential tensor onto the spherical harmonic tensors, a set of nine 3 $\times$ 3 matrices whose complex-valued expansion coefficients can be directly interpreted in terms of phase, absorption, birefringence, and diattenuation. We schematize each of these spherical harmonic tensors in Ext. Data Fig. 1a by drawing each tensor’s eigenvalues and eigenvectors, and we describe the spherical harmonic tensor basis in detail in Supp. 3.
@@
In a fluorescent sample, the 3D emission dipole moment can be represented by a three-component vector $f^{(\text{fl})}_{i}$ (Fig. 1a, vi; Ext. Data Fig. 1b) with real-valued coefficients for purely linear dipoles and complex-valued coefficients for arbitrary dipoles. Contrast arises from the scatter-scatter term, so our measurements are proportional to the squares of the dipole components $|f^{(\text{fl})}_{i}|^{2}$.
@@
For dynamic ensembles of fluorescent emitters, the measurements are proportional to the second-moment matrix $\langle f^{(\text{fl})}_{i}f^{*(\text{fl})}_{j}\rangle$ [backer_extending_2014, zhang_six-dimensional_2023, zhang_single-molecule_2021]. Similar to the scattering potential tensor, we can expand the second-moment matrix onto the spherical harmonic tensors, but here we interpret the coefficients in terms of orientation distribution functions [chandler_spatio-angular_2019, chandler_volumetric_2025].
@@
Finally, we express our data in terms of the Stokes parameters, a set of four real-valued parameters that are physically interpretable as the intensities measured behind various polarizing filters. The Stokes parameters can also be interpreted as the coefficients of the electric field’s second-moment matrix expanded onto the Pauli matrices (Ext. Data Fig. 1c, Supp. 4).
vector scalar
Table 1: WaveOrder’s transfer functions in terms of autocorrelated pupils. Transfer functions for different contrast modes (rows) and optical models (columns) expressed in terms of the illumination model $S$, the scattering model $G$, the detection model $P$, the object-space basis $\mathcal{Y}$, and the data-space basis $\sigma$. Sums over $i,i^{\prime},j,j^{\prime},k$ and $k^{\prime}$ are implied.
4.2.6 Contrast-separable imaging systems
In many microscopy imaging systems, data in each channel is the sum of the contributions from each material property. We call these systems channel linear and we can express the forward operator as
$\mathbf{d}_{c}=\mathcal{H}_{cp}\mathbf{f}_{p}+\mathbf{b}_{c}$ (8)
where $c$ indexes channels and $p$ indexes material properties. For example, multi-channel fluorescence microscopy often suffers from crosstalk, but these systems are still channel linear if the data is the sum of the contribution from each type of fluorophore.
When there is no crosstalk between channels (for example, when fluorescence filters are perfect), we say the imaging system is channel separable, which implies that the system can be written as
$\mathbf{d}_{c}=\mathcal{H}_{cc}\mathbf{f}_{c}+\mathbf{b}_{c},$ (9)
i.e. the operator $\mathcal{H}$ is diagonal over channels.
The WaveOrder framework considers imaging systems where some groups of channels are separable (e.g. fluorescence contrast), and some groups of channels are merely linear (e.g. label-free contrast). We call such imaging systems contrast separable and they can be expressed as
$\mathbf{d}^{(m)}_{c}=\mathcal{H}^{(m)}_{cp}\mathbf{f}^{(m)}_{p}+\mathbf{b}_{c}^{\text{(}m)},$ (10)
where $m$ indexes each of the $M$ contrast modes. For example, an imaging system with two channel-separable fluorescent channels and four channel-linear label-free data channels that jointly encode three object properties can be written as
$\begin{bmatrix}\mathbf{d}_{1}^{(1)}\\ \mathbf{d}_{1}^{(2)}\\ \mathbf{d}_{1}^{(3)}\\ \mathbf{d}_{2}^{(3)}\\ \mathbf{d}_{2}^{(3)}\\ \mathbf{d}_{3}^{(3)}\end{bmatrix}=\begin{bmatrix}\mathcal{H}^{(1)}_{1,1}&0&0&0&0\\ 0&\mathcal{H}^{(2)}_{1,1}&0&0&0\\ 0&0&\mathcal{H}^{(3)}_{1,1}&\mathcal{H}^{(3)}_{1,2}&\mathcal{H}^{(3)}_{1,3}\\ 0&0&\mathcal{H}^{(3)}_{2,1}&\mathcal{H}^{(3)}_{2,2}&\mathcal{H}^{(3)}_{2,3}\\ 0&0&\mathcal{H}^{(3)}_{3,1}&\mathcal{H}^{(3)}_{3,2}&\mathcal{H}^{(3)}_{3,3}\\ 0&0&\mathcal{H}^{(3)}_{4,1}&\mathcal{H}^{(3)}_{4,2}&\mathcal{H}^{(3)}_{4,3}\end{bmatrix}\begin{bmatrix}\mathbf{f}^{(1)}_{1}\\ \mathbf{f}^{(2)}_{1}\\ \mathbf{f}^{(3)}_{1}\\ \mathbf{f}^{(3)}_{2}\\ \mathbf{f}^{(3)}_{3}\end{bmatrix}+\begin{bmatrix}\mathbf{b}_{1}^{(1)}\\ \mathbf{b}_{1}^{(2)}\\ \mathbf{b}_{1}^{(3)}\\ \mathbf{b}_{2}^{(3)}\\ \mathbf{b}_{2}^{(3)}\\ \mathbf{b}_{3}^{(3)}\end{bmatrix}$. (11)
In other words, we assume that the forward operator is block diagonal over channels, so we can split our reconstruction problem into subproblems, one for each contrast mode
$\mathbf{\hat{f}}_{p}^{(m)}=\mathcal{R}_{\eta}^{(m)}\mathbf{d}_{c}^{(m)}=\operatorname*{argmin}_{\mathbf{f}_{p}^{(m)}}Q_{\eta}^{\text{(tls)}}\left(\mathbf{f}_{p}^{(m)},\mathbf{d}_{c}^{(m)}\right),$ (12)
where we have defined $\mathcal{R}_{\eta,pc}^{(m)}$ as a reconstruction operator for each contrast mode. Eq. 12 is the core of our reconstruction algorithm.
4.2.7 Linear contrast-separable shift-invariant imaging systems
@@
When a contrast-separable imaging system is approximately spatially linear and shift invariant, we can express our imaging model as
$d^{(m)}_{c}(\mathbf{r}_{d})=\sum_{p}\int d\mathbf{r}_{o}\,h^{(m)}_{cp}(\mathbf{r}_{d}-\mathbf{r}_{o})f^{(m)}_{p}(\mathbf{r}_{o})+b_{c}^{(m)},$ (13)
@@
where $m$ indexes contrast mmodes, $p$ indexes material pproperties, $c$ indexes data cchannels, and $h^{(m)}_{cp}$ is a bank of point spread functions that model the entire multi-contrast multi-channel imaging system. We can reexpress this relationship in the Fourier domain as
@@
4.2.8 Summary of transfer functions
@@
The Green’s tensor spectrum is particularly important for modeling anisotropic contrast. Linear dipole moments emit polarized light in a doughnut-shaped intensity pattern (Ext. Data Fig. 2a), and the Green’s tensor spectrum (Ext. Data Fig. 2b) efficiently models all dipole emitters (coherent or incoherent; linear, circular, or elliptical dipoles in any orientation) with a single function.
All transfer functions can be expressed as products and autocorrelations of the illumination, scattering, and detection models, see Table 1 and Supp. 5. We refer to the complete transfer functions as vector models because they account for the complete vectorial nature of light and dipole scattering, and we also include scalar models that ignore vector effects, which are reasonable approximations when unpolarized illumination and unpolarized detection are used on isotropic samples.
Ext. Data Figs. 2c–f show the support and phase of several examples of WaveOrder’s transfer functions. We briefly highlight several key features